Larry invested part of his $38,000 advance at 6% annual simple interest and the rest at 9% annual simple interest. If his total yearly interest from both accounts was $2,850, find the amount invested at each rate. At 6% and 9%.
Answer by mananth(16946)

Part I 6.00% per annum ------------- x
Part II 9.00% per annum ------------ y
38000
Interest----- 2850

Part I 6.00% per annum ---x
Part II 9.00% per annum ---y
Sum of investments
x + 1 y= 38000 -------------1
Interest earned on both investments
6.00% x + 9.00% y= 2850
Multiply by 100
6 x + 9 y= 285000.00 --------2
Multiply (1) by -6
we get
-6 x -6 y= -228000.00
Add this to (2)
0 x 3 y= 57000
divide by 3
y = 19000
Part I 9.00% $ 19000
Part II 6.00% $ 19000

CHECK
19000 --------- 9.00% ------- 1710.00
19000 ------------- 6.00% ------- 1140.00
Total -------------------- 2850.00

m.ananth@hotmail.ca